OpenStudy (anonymous):

what is 1.05 as a percent

OpenStudy (solomonzelman):

multiply 1.05 times 100.

OpenStudy (solomonzelman):

this will get you 1.05 in terms of a percentage.

OpenStudy (anonymous):

Remove the decimal point, and you get your answer.

OpenStudy (solomonzelman):

in this case, it is so... :D

OpenStudy (anonymous):

1.05 As a fraction: 1.05 = 1 5/100 = 1 1/20 As a percent: 1.05 x 100 =105%. OR 1 1/20 x 100 =21/20 x 100 = (21x100) ÷20 = 2100 ÷20=105%